OpenStudy (anonymous):

Linear Systems in Two Variables with applications. solve using elimination or substitution. 4x + 3/4y = 14 -9x+ 5/8y = -13

OpenStudy (anonymous):

Do you have an idea on what would be easier to use, elimination or substitution?

OpenStudy (anonymous):

not really... i know elimination mostly.

OpenStudy (anonymous):

Alright, well lets go with elimination then. (For this case it appears to be the easier way to solve it)

OpenStudy (anonymous):

What variable do you want to get rid of first?

OpenStudy (jennychan12):

\[4x+\frac{ 3y }{ 4 } = 14\] \[-9x +\frac{ 5y }{ 8 } = -13\] to make it easier, multiply each equation by the denominator \[16x+ 3y = 56 \] call this equation 1 \[-72x +5y = -104\] call this equation 2 the next few steps requires algebra. ELIMINATION multiply equation 1 by -5 and equation 2 by 3 \[-80x- 15y = -280 \] \[-216x +15y = -312\] add \[-296x = 592 \] \[x = 2 \] plug this back into either equation 1 or 2 I'll choose equation 1. \[16(2)+ 3y = 56 \] \[y=3\]

OpenStudy (anonymous):

that is perfect.. thank you so much, i didn't realize i could mulitply everything by the denominator.

OpenStudy (jennychan12):

yeah, sometimes, you just hafta learn certain tricks to stuff. cuz you could've done it with fractions, but those are a nightmare....